OpenStudy (anonymous):

Solve. (x – 5)^2 = 2 Using quadratic formula

OpenStudy (earthcitizen):

Quadratic formula state\[x = (-b \pm \sqrt{b^{2}-4ac})/2a\] Expand (-5+x)^2 = 2 (-5+x) (-5+x) = 2 -5 (-5+x)+x (-5+x) = 2 -5 (-5+x)+x (-5)+x x = 2 -5 (-5+x)-5 x+x x = 2 -5 (-5)-5 x-5 x+x x = 2 -5 (-5)+2 (-5) x+x x = 2 -5 (-5)-5 2 x+x x = 2 (-5)^(1+1)-5 2 x+x^(1+1) = 2 25-5 2 x+x^2 = 2 25-10 x+x^2 = 2 Rearrange, \[x^2-10 x+23 = 0\] from the equation above a =1, b =-10 and c =23, plug into the formula chug! chug!

OpenStudy (anonymous):

whats the value inside the radical

OpenStudy (earthcitizen):

kk, equate \[ax^2 +bx+c=x^2 -10x+23=0\] and solve for x from the formula